# DreamControl: Human-Inspired Whole-Body Humanoid Control for Scene Interaction via Guided Diffusion

### v1

@@ IX-A Model architecture
IX-B Environment parameters
We randomize the object/target location (by sampling different trajectories in some tasks), object mass, friction of surface/object. All of these variations for each task are listed in Table XI.
TABLE XI: Environment randomization parameters for each task
@@ IX-B Environment parameters
IX-C Task-specific sparse rewards
We add task-specific sparse rewards that are chosen to be indicative of whether the task is successful as well. These rewards are listed in Table XII. For each task, we have a body part link, $b$ that maybe used in the definition of the sparse rewards and the time $t_{p}$ when the interaction is supposed to happen.
IX-D Task-specific dense rewards for TaskOnly+
@@
$r_{t}^{\text{dense}}=\lVert p_{t-\Delta t}^{\text{b}}-p_{t_{p}}^{\text{ref,b}}\rVert_{2}-\lVert p_{t}^{\text{b}}-p_{t_{p}}^{\text{ref,b}}\rVert_{2}$.
The specific body parts used in each task are listed in Table XII.
TABLE XII: Task-specific sparse rewards for each task
@@
IX-E Reward weights for each task
The rewards weights for each task are listed in Table XIII.
TABLE XIII: Reward weights for all tasks

### v2

@@ X-A Model architecture
X-B Environment parameters
We randomize the object/target location (by sampling different trajectories in some tasks), object mass, friction of surface/object. All of these variations for each task are listed in Table XII.
TABLE XII: Environment randomization parameters for each task
@@ X-B Environment parameters
X-C Task-specific sparse rewards
We add task-specific sparse rewards that are chosen to be indicative of whether the task is successful as well. These rewards are listed in Table XIII. For each task, we have a body part link, $b$ that maybe used in the definition of the sparse rewards and the time $t_{g}$ when the interaction is supposed to happen.
X-D Task-specific dense rewards for TaskOnly+
@@
$r_{t}^{\text{dense}}=\lVert p_{t-\Delta t}^{\text{b}}-p_{t_{g}}^{\text{ref,b}}\rVert_{2}-\lVert p_{t}^{\text{b}}-p_{t_{g}}^{\text{ref,b}}\rVert_{2}$.
The specific body parts used in each task are listed in Table XIII.
TABLE XIII: Task-specific sparse rewards for each task
@@
X-E Reward weights for each task
The rewards weights for each task are listed in Table XIV.
TABLE XIV: Reward weights for all tasks